3x-3=3y-6 =z-7,find the unit vector parallel to the above line

1
Expert's answer
2020-08-06T17:42:50-0400

Let us rewrite the equation in form

"\dfrac{x-1}{\frac{1}{3}} = \dfrac{y-2}{\frac{1}{3}} = \dfrac{z-7}{1}" .

In such form the denominators represent the coordinates of the vector parallel to the line. The coordinates are (1/3, 1/3, 1). But we can see the vector is not unit, so we should normalize it.

The length of the vector above is

"\sqrt{(1/3)^2+(1/3)^2+1^2} = \dfrac{\sqrt{11}}{3}."

Therefore, the unit vector will have coordinates "\left(\dfrac{1/3}{\sqrt{11}/3}\,,\dfrac{1/3}{\sqrt{11}/3}\,, \dfrac{1}{\sqrt{11}/3}\right)" , after simplifying

"\left( \dfrac{\sqrt{11}}{11}\,, \dfrac{\sqrt{11}}{11}\,, \dfrac{3\sqrt{11}}{11}\right)" .